is the value of sin(3t-540)can be written as -sin3t?

Answer by stanbon(57967)

is the value of sin(3t-540)can be written as -sin3t?
-----
Note: sin(-x) = -sin(x)
cos(-x) = cos(x)
------------------------------
sin(3t-540) = sin(3t)*cos(-540)+sin(-540)(cos(3t)
---
= sin(3t)(cos(540))+[-sin(540)(cos(3t)
---
= sin(3t5)(-1)+[-0*cos(3t)]
---
= -sin(3t)
===============
Cheers,
Stan H.
===============